OpenStudy (anonymous):

4^x= 1÷64 how would you solve that

OpenStudy (caozeyuan):

\[4^{x}=\frac{ 1 }{ 64 }\]

OpenStudy (caozeyuan):

\[\ln 4^{x}=\ln 64^{-1}\]

OpenStudy (caozeyuan):

\[x \ln 4=-\ln 64\]

OpenStudy (caozeyuan):

\[x=-\frac{ \ln 64 }{ \ln 4 }\]

OpenStudy (caozeyuan):

the rat can be done by calculator